# I’M HOI: Inertia-aware Monocular Capture of 3D Human-Object Interactions

### v1

@@ 4 Dataset
[…]
@@
Given multi-view videos, we reproduced human motions in SMPL-H format [68] using an open-source toolbox [1]. In order to track accurate object pose in 3D scene, we first manually annotated single key-frame segmentation in all views, and broadcasted to the whole sequence [90, 35, 8, 45]. Subsequently, we optimized euclidean transformations, which can exactly define object motions, by fitting reprojected silhouette to multi-view masks. For object geometries, we utilized a public application [60] to get 3D scan templates. In terms of inertial signals, we adopted primitive rotation data $\bm{R}_{s}$ in matrix form and transformed raw acceleration $\bm{a}_{raw}$ in sensor coordinate to free acceleration $\bm{a}_{free}$ in global coordinate through $\bm{a}_{free}=\bm{R}_{s}\bm{a}_{raw}-\bm{g}$, where $\bm{g}=[0,0,9.81]^{\mathrm{T}}$ is the gravitational acceleration.
Figure 5: Qualitative comparison results. I’m-HOI performs better than baselines and generalizes well to new datasets.
@@ 5 Experiments
In this section, we first introduce the datasets and metrics used for training and evaluation. Then we comprehensively compare our approach to baseline methods. We also extensively conduct ablation studies to demonstrate the effectiveness of pivot components in our network design and the necessity of IMU modality.
@@ 5.1 Datasets and Evaluation Metrics
[…]
@@
  • Sliding Window Chamfer Distance ($cm$) [86] computes the chamfer distance in the same way but performing holistic Procrustes alignment on combined mesh of 10-second results with the ground truth.
@@ Results.
In order to validate the generalization ability, we apply purely optimization-based method PHOSA [104], learning-and-optimization methods CHORE[85], VisTracker[86] and our approach trained on BEHAVE[4] and InterCap[24] dataset to HODome[102] and CHAIRS[26] dataset. As shown in Table 2, I’m-HOI generalizes better than baselines by a large margin and achieves more balanced performance between per-frame and sequential results. The illustration of Figure 5 shows that I’m-HOI can adapt to various scene backgrounds and objects.
@@
[…]
@@
Table 2: Quantitative evaluations of generalization ability. I’m-HOI outperforms baselines by a large margin on new data.
5.3 Ablation Study
[…]
Network Architecture.
Table 3 shows the performance of models with or without the mesh alignment feedback (maf.), optimization module (optim.) and diffusion filter (filter.). It is demonstrated that maf. improves per-frame object tracking results and optim. brings better temporal consistency. Moreover, filter. further corrects human-object spatial arrangements onto the learned interaction manifold. Compared with naive implementation, the full pipeline of I’m-HOI performs 4 times better. Figure 6 illustrates that inaccurate prediction is corrected progressively when maf. and optim. are applied. Also, the hand motion generated from filter. makes the capture result more vivid and realistic.
@@ 5.3 Ablation Study
Table 3: Quantitative evaluations of network architecture design.
@@
We experiment on several different baselines with IMU modality input by adding extra inertial optimization term described in Equation 4 to their pipelines. Qualitative results illustrated in Figure 7 clearly exhibit the improvements on object pose estimation of baselines after introducing IMU modality, compared with Figure 5. Quantitative performance reported in Table 4 shows decent increase of baselines’ performance, compared with the statistics in Table 1. Additionally, we observe that our approach achieves better results when the IMU modality is involved, especially for object tracking. Moreover, Table 4 indicates that naively incorporating IMU modality input within baselines is unable to maximize its advantages, which further verifies the effectiveness of our network design.
@@ 5.4 Limitation
As the first trial to explore challenging 3D human-object interactions capture using a minimal amount of RGB camera and object-mounted IMU sensor, the proposed I’m-HOI still owns limitations as follows. First, our method relies on pre-scanned object template and manual manipulation of sensor-template coordinate alignment. Our method is also restricted to rigid object tracking. It’s a promising direction to extend the interesting task to articulated or even deformable objects in template-free framework.
@@ 6 Conclusion
[…]
@@ A.1 General Interaction Motion Inference
We use the pre-trained ResNet-34 [17] $f^{\text{enc}}:\mathbb{R}^{H\times W\times 4}\mapsto\mathbb{R}^{8\times 8\times 512}$ to extract image feature, where $H=W=256$. Then we use 3 stacked deconvolution layers to construct feature pyramid, each of which $f^{\text{deconv}}_{i}:\mathbb{R}^{H_{\text{in}}\times W_{\text{in}}\times C_{\text{in}}}\mapsto\mathbb{R}^{2H_{\text{in}}\times 2W_{\text{in}}\times C_{\text{out}}}$ takes input feature with resolution $H_{\text{in}}=W_{\text{in}}=8,16,32$ and channel $C_{\text{in}}=512,256,256$ respectively, and outputs 256-channel 2 times upsampled feature map. Each deconvolution layer is followed by a Batch Normalization layer and ReLu activation function. For every single intermediate feature map, one regressor $f^{\text{reg}}_{i}:\mathbb{R}^{H_{\text{in}}\times W_{\text{in}}\times 256}\mapsto\mathbb{R}^{6+3}$ is tailored to embed mesh-aligned feature to $\mathbb{R}^{2000}$ and concatenates it with $\hat{\bm{R}}_{o}^{(i-1)},\hat{\bm{T}}_{o}^{(i-1)}$ to predict $\Delta\hat{\bm{R}}_{o}^{(i)},\Delta\hat{\bm{T}}_{o}^{(i)}$. All regressors consist of two hidden Linear layers with dimension $1024$ and two output Linear layers predicting delta rotation and translation separately. Dropout with $0.5$ probability are inserted between each two adjacent Linear layers.
@@
The proposed $f^{\text{enc}},\{f^{\text{deconv}}_{i}\}_{i=1}^{3},\{f^{\text{reg}}_{i}\}_{i=1}^{3}$ are trained end-to-end with the inverse kinematics layer, supervised by $\mathcal{L}=\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}+\mathcal{L}_{\text{twist}}+\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$. Particularly, the object-oriented mesh alignment feedback loss $\mathcal{L}_{\text{maf}}=\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$ is added after 55 training epochs. The loss weights are: $\lambda_{\text{j2d}}=1\times 10^{-9},\lambda_{\text{occ-sil}}=1\times 10^{-6},\lambda_{\text{area}}=2\times 10^{-7}$. The model is trained for 190 epochs on 6 NVIDIA GeForce RTX 3090 GPUs. In each epoch, we randomly downsample images 8$\times$ to train and the training batch size is set to 8.
@@
[…]
@@ A.2 Category-specific Motion Diffusion Filter
We use 4 transformer encoder-only layers with 4 attention heads to learn category-specific human-object interaction manifold. The model dimension $D_{\text{model}}=1024$ and key, value dimension $D_{\text{key}}=D_{\text{value}}=512$. We take standard $N=1000$ denoising steps and sinusoidal positional encoding function. Different from [43] in which the condition is exactly a part of the target motion, we take results from previous stage together with raw IMU measurements as condition to model the transformation from prediction distribution to the true manifold.
@@
We first warm up the diffusion model on our whole training dataset only with simple objective function for 100 epochs. After that, we train the model on category-specific data with designed regularization terms $\mathcal{L}_{\text{consist}},\mathcal{L}_{\text{vel}}$ and $\mathcal{L}_{\text{imu}}$ to implicitly model particular interaction patterns. The regularization term weights are $\lambda_{\text{off}}=\lambda_{\text{vel}}=\lambda_{\text{consist}}=1,\lambda_{\text{imu}}=100$. More detailed, we apply $\mathcal{L}_{\text{off}}$ and $\mathcal{L}_{\text{vel}}$ for 35 epochs before adding $\mathcal{L}_{\text{consist}}$ and $\mathcal{L}_{\text{imu}}$. For better generation, we maintain an exponential moving average (EMA) version model during training, which is updated per 10 epochs with decay rate $0.995$. We also enable automatic mixed precision (AMP) to accelerate training process. The model is trained for 55 epochs on single NVIDIA GeForce RTX 3090 GPU. The training batch size is set to 128.
@@
[…]
@@ Appendix B Data Preparation Details
To synchronize RGB data and IMU measurements, we asked the performer to wear an extra IMU sensor on his/her ankle part and make a takeoff move at the beginning of each interaction split. We automatically identified the moment the performer falling to the ground through gravitational acceleration mutation from IMU signals, and manually annotated that in RGB sequences.
@@
To mitigate spatial misalignment between camera and IMU, we conducted spatial alignment once per ten minutes. Specifically, in our multi-modal multi-sensor system, there exists multiple coordinate frames, including $\{\mathcal{F}_{C_{i}}\}_{i=0}^{31}$ for cameras, $\mathcal{F}_{W}$ for world and $\mathcal{F}_{I}$ for inertia. Since the transformation $\mathcal{T}_{W\rightarrow C_{i}}\in\mathcal{SO}(3)$ from $\mathcal{F}_{W}$ to $\mathcal{F}_{C_{i}}$ is easy to obtain through off-the-shelf toolbox, our goal is to calibrate the transformation $\mathcal{T}_{I\rightarrow W}\in\mathcal{SO}(3)$ from $\mathcal{F}_{I}$ to $\mathcal{F}_{W}$.
In implementation, we capture the global orientation $\{\bm{R}_{t}^{W}\in\mathcal{SO}(3)\}_{t=0}^{T-1}$ of a performer who wears an IMU sensor on his/her waist part and circles around in $\mathcal{F}_{W}$ using [1], and record the inertial rotation measurement $\{\bm{R}_{t}^{I}\in\mathcal{SO}(3)\}_{t=0}^{T-1}$ in $\mathcal{F}_{I}$. Suppose that the IMU sensor is relatively fixed to the performer, we can construct the following equation:
@@ B.2 Acceleration Data Processing
Assume that all the objects are rigid and have uniform rotational inertia, the centroid of each object is equivalent to its geometry center. However, it’s impractical to mount IMU sensor exactly on the center that is probably inside of the object. Hence, extra linear acceleration can be generated even from purely rotational motion, which is undesirable. To eliminate such noises, we first fix a mounting point for each object and manually measure the directional offset $\vec{\bm{r}}$ from center to that point in mesh processing software [9].
@@
According to recorded angular velocity $\vec{\bm{w}}$, the extra linear velocity generated from rotation is $\vec{\bm{v}}=\vec{\bm{w}}\times\vec{\bm{r}}$, and we can determine the extra linear acceleration: $\delta\bm{a}_{t}=\frac{\bm{v}_{t}-\bm{v}_{t-\Delta t}}{\Delta t}$. Finally, the normalized acceleration data can be achieved by subtracting $\delta\bm{a}_{t}$ from the raw measurements.
@@
We also simulate synthetic IMU data based on ground-truth object motion annotations of [4, 24, 102, 26]. In particular, for inertial acceleration data, we follow [93, 46, 67] to calculate second-order difference of object translation:
@@
where $n=4$ is the smoothing factor for better approximation to real acceleration, and $\tau=\frac{1}{\text{fps}}$ is the time interval between two consecutive frames.

### v2

@@ 4 Dataset
To train and evaluate our I’m-HOI, we collect an Inertial and Multi-view Highly Dynamic human-object interactions Dataset (IMHD2), consisting of human, object motions, inertial measurements and object 3D scans.
@@
Given multi-view videos, we reproduced human motions in SMPL-H format [68] using an open-source toolbox [1]. To accurately track object pose in a 3D scene, we manually annotated single key-frame segmentation in all views and broadcasted it to the entire sequence [90, 36, 8, 46]. Subsequently, we optimized Euclidean transformations, which precisely define object motions, by fitting reprojected silhouette to multi-view masks. For object geometries, we utilized a public application [61] to obtain 3D scan templates. In terms of inertial signals, we adopted primitive rotation data $\bm{R}_{s}$ in matrix form and transformed raw acceleration $\bm{a}_{raw}$ in sensor coordinate to free acceleration $\bm{a}_{free}$ in global coordinate through $\bm{a}_{free}=\bm{R}_{s}\bm{a}_{raw}-\bm{g}$, where $\bm{g}=[0,0,9.81]^{\mathrm{T}}$ is the gravitational acceleration.
Figure 5: Qualitative comparison results. I’m-HOI outperforms the baselines and generalizes well to new datasets.
@@ 5 Experiments
In this section, we first introduce the datasets and metrics used for training and evaluation. We then provide a comprehensive comparison between our approach with baseline methods. We also perform extensive ablation studies to demonstrate the effectiveness of pivot components in our network design and the necessity of the IMU modality.
@@ 5.1 Datasets and Evaluation Metrics
We train I’m-HOI using BEHAVE [4], InterCap [24] and IMHD2, and evaluate it on five datasets which also include HODome [102] and CHAIRS [26]. We adhere to the official train-test data partitioning of BEHAVE and InterCap, which is established by VisTracker [86]. Given the relatively slow inference speeds of baselines [104, 85, 86], we curate partial yet representative data from IMHD2, HODome, and CHAIRS to construct sub test sets for thorough evaluation.
@@
  • Sliding Window Chamfer Distance ($cm$) [86] computes the chamfer distance in the same way but performing holistic Procrustes alignment on the combined mesh of 10-second results with the ground truth.
@@ Results.
To assess the generalization capabilities, we evaluate the performance of purely optimization-based method PHOSA [104], learning-and-optimization methods CHORE[85] and VisTracker[86], as well as our proposed approach trained on BEHAVE[4] and InterCap[24], across HODome[102] and CHAIRS[26]. As shown in Table 2, I’m-HOI generalizes better than the baselines by a large margin and achieves more balanced performance between per-frame and sequential results. Furthermore, Figure 5 demonstrates the adaptability to diverse scenarios of I’m-HOI.
@@
We conduct a comparative analysis of the inference efficiency across different methods using a specific sequence from InterCap dataset [24]. Among the methods evaluated, the purely optimization-based framework PHOSA [104] takes the longest inference time which is approximately 2 minutes per frame. CHORE [85] speeds up to 1 minute per frame, while VisTracker [86] further reduces the time cost to 20 seconds. Notably, I’m-HOI requires only about 0.5 seconds per frame for the complete pipeline. It is worth mentioning that omitting the optional optimization module could lead to additional enhancements in efficiency.
5.3 Ablation Study
Extensive ablation studies are conducted on IMHD2 to evaluate our network architecture design and IMU modality.
Network Architecture.
Table 3 shows the performance of models with and without the mesh alignment feedback (maf.), optimization module (optim.) and diffusion filter (filter.). It is demonstrated that maf. improves per-frame object tracking results and optim. brings better temporal consistency. In addition, filter. further corrects human-object spatial arrangements onto the learned interaction manifold. Compared to the naive implementation, the full pipeline of I’m-HOI performs 4 times better. Figure 6 illustrates that inaccurate prediction is progressively corrected when maf. and optim. are applied. Also, the hand motion generated by filter. makes the capture result more vivid and realistic.
@@ 5.3 Ablation Study
Table 2: Quantitative evaluations of generalization ability.
@@
Table 3: Quantitative evaluation of network architecture design.
@@
We experiment on several different baselines with IMU modality input by adding an additional inertial optimization term described in Equation 4 to their pipelines. The qualitative results shown in Figure 7 clearly demonstrate the improvements in object pose estimation of the baselines after introducing the IMU modality, compared to Figure 5. The quantitative performance reported in Table 4 shows a decent increase in the performance of baselines, compared to the statistics in Table 1. Additionally, we observe that our approach achieves better results when the IMU modality is involved, especially for object tracking. Furthermore, Table 4 shows that naively incorporating the IMU modality input into baselines is unable to maximize its benefits, which further verifies the effectiveness of our network design.
@@ 5.4 Limitation
The proposed I’m-HOI is the first trial to explore challenging 3D human-object interactions capture using a minimal amount of RGB camera and object-mounted IMU sensor. However, it still has limitations. Firstly, our method relies on pre-scanned object templates and manual manipulation of sensor-template coordinate alignment. Additionally, our method is restricted to rigid object tracking. Extending this method to articulated or even deformable objects in a template-free framework is promising.
@@ 6 Conclusion
Acknowledgement.
This work was supported by National Key R&D Program of China (2022YFF0902301), Shanghai Local college capacity building program (22010502800). We also acknowledge support from Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI).
@@ Acknowledgement.
In this supplementary, we commence by providing a comprehensive exposition on the implementation details of our method. Subsequently, we expound upon the calibration of the multi-modal system and the processing of IMU acceleration data during the data collection phase of IMHD2. Finally, we present additional qualitative and quantitative results to further validate the efficacy of I’m-HOI, with a particular focus on the evaluation of each regularization term integrated within the category-specific interaction diffusion filter component.
@@ A.1 General Interaction Motion Inference
We employ a pre-trained ResNet-34 model [17] $f^{\text{enc}}:\mathbb{R}^{H\times W\times 4}\mapsto\mathbb{R}^{8\times 8\times 512}$ to extract image features, where $H=W=256$. Subsequently, we utilize 3 stacked deconvolution layers to construct a feature pyramid, with each layer $f^{\text{deconv}}_{i}:\mathbb{R}^{H_{\text{in}}\times W_{\text{in}}\times C_{\text{in}}}\mapsto\mathbb{R}^{2H_{\text{in}}\times 2W_{\text{in}}\times C_{\text{out}}}$ receiving input features with resolution $H_{\text{in}}=W_{\text{in}}=8,16,32$ and channel $C_{\text{in}}=512,256,256$ respectively, and producing a 256-channel feature map that is upsampled by a fator of 2. Following each deconvolution layer are Batch Normalization and ReLu activation layers. For every intermediate feature map, a specific regressor $f^{\text{reg}}_{i}:\mathbb{R}^{H_{\text{in}}\times W_{\text{in}}\times 256}\mapsto\mathbb{R}^{6+3}$ is tailored to embed it to $\mathbb{R}^{2000}$ and concatenates it with $\hat{\bm{R}}_{o}^{(i-1)},\hat{\bm{T}}_{o}^{(i-1)}$ to predict $\Delta\hat{\bm{R}}_{o}^{(i)},\Delta\hat{\bm{T}}_{o}^{(i)}$. Each regressor comprises two hidden Linear layers with a dimension of $1024$, as well as two output Linear layers that predict delta rotation and translation independently. Dropout layers with a probability of $0.5$ are inserted between each pair of consecutive Linear layers.
@@
The proposed $f^{\text{enc}},\{f^{\text{deconv}}_{i}\}_{i=1}^{3},\{f^{\text{reg}}_{i}\}_{i=1}^{3}$ are trained end-to-end with the inverse kinematics layer, supervised by $\mathcal{L}=\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}+\mathcal{L}_{\text{twist}}+\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$. Particularly, the object-oriented mesh alignment feedback loss $\mathcal{L}_{\text{maf}}=\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$ is added after 55 training epochs. The loss weights are: $\lambda_{\text{j2d}}=1\times 10^{-9},\lambda_{\text{occ-sil}}=1\times 10^{-6},\lambda_{\text{area}}=2\times 10^{-7}$. The model is trained for 190 epochs on 6 NVIDIA GeForce RTX 3090 GPUs. In each epoch, we randomly sample one from 8 images to train. The training batch size is set to 8.
@@
The optimization energy function defined as $\mathcal{E}=w_{\text{visual}}\mathcal{E}_{\text{visual}}+w_{\text{imu}}\mathcal{E}_{\text{imu}}$ is configured with: $w_{\text{visual}}=20,w_{\text{imu}}=1\times 10^{5}$. We set the learning rate during optimization to $0.01$ for 30-fps data (BEHAVE [4], InterCap [24] and CHAIRS [26]) and $5\times 10^{-4}$ for 60-fps data (IMHD2 and HODome [102]).
@@ A.2 Category-specific Motion Diffusion Filter
We employ 4 transformer encoder-only layers, each equipped with 4 attention heads, to learn category-specific human-object interaction manifold. The model dimension $D_{\text{model}}=1024$ and the key, value dimension $D_{\text{key}}=D_{\text{value}}=512$. We take $N=1000$ steps and sinusoidal positional encoding function during denoising phase. In contrast to the methodology outlined in [44], where the condition is exactly a part of the target motion, we leverage outcomes from the preceding stage alongside raw IMU measurements as conditions to model the transition from the predictive distribution to the authentic manifold.
@@
We initially warm up the diffusion model solely on our complete training dataset using simple objective function for 100 epochs. Subsequently, we proceed to train the model on category-specific data, incorporating specially designed regularization terms $\mathcal{L}_{\text{consist}},\mathcal{L}_{\text{vel}}$ and $\mathcal{L}_{\text{imu}}$ to implicitly model distinct interaction patterns. The regularization term weights are $\lambda_{\text{off}}=\lambda_{\text{vel}}=\lambda_{\text{consist}}=1,\lambda_{\text{imu}}=100$. More detailed, we apply $\mathcal{L}_{\text{off}}$ and $\mathcal{L}_{\text{vel}}$ for 35 epochs before adding $\mathcal{L}_{\text{consist}}$ and $\mathcal{L}_{\text{imu}}$. To enhance the generation results, we maintain an exponential moving average (EMA) version of the model throughout training, updating it every 10 epochs with a decay rate of $0.995$. Additionally, we leverage Automatic Mixed Precision (AMP) to accelerate the training procedure. The model is trained for 55 epochs on a single NVIDIA GeForce RTX 3090 GPU, with the training batch size set to 128.
@@
Table 5: IMHD2 collects 10 distinct objects along with a range of interaction motions associated with each object.
@@ B.1 System Calibration of IMHD2
In order to synchronize RGB data with IMU measurements, we instructed the performer to wear an additional IMU sensor on the ankle area and execute a takeoff motion at the onset of each interaction segment. By detecting the point at which the performer falls to the ground based on the gravitational acceleration mutation in the IMU signals, we automatically pinpointed this moment as the starting frame and manually annotated it within the RGB sequences.
@@
To mitigate spatial misalignment between camera and IMU, we conducted spatial alignment once per ten minutes. Specifically, in our multi-modal and multi-sensor system, there exists multiple coordinate frames, including $\{\mathcal{F}_{C_{i}}\}_{i=0}^{31}$ for cameras, $\mathcal{F}_{W}$ for world and $\mathcal{F}_{I}$ for inertia. Since the transformation $\mathcal{T}_{W\rightarrow C_{i}}\in\mathcal{SO}(3)$ from $\mathcal{F}_{W}$ to $\mathcal{F}_{C_{i}}$ is easy to obtain through off-the-shelf multi-camera calibration toolbox, our goal is to calibrate the transformation $\mathcal{T}_{I\rightarrow W}\in\mathcal{SO}(3)$ from $\mathcal{F}_{I}$ to $\mathcal{F}_{W}$.
In our implementation, we capture the global orientation $\{\bm{R}_{t}^{W}\in\mathcal{SO}(3)\}_{t=0}^{T-1}$ of the performer who circles around in $\mathcal{F}_{W}$ by [1]. The inertial rotation measurement $\{\bm{R}_{t}^{I}\in\mathcal{SO}(3)\}_{t=0}^{T-1}$ in $\mathcal{F}_{I}$ is simultaneously recorded by an IMU sensor positioned at the waist area. Suppose the IMU sensor is relatively fixed to the performer, we can construct the following equation:
@@ B.2 Acceleration Data Processing
Given the assumption that all objects are rigid and possess uniform rotational inertia,
@@
with their centroids equivalent to their geometry centers, practical constraints arise when attempting to mount the IMU sensor precisely onto these centers. which may lie within the object. Consequently, extraneous linear acceleration may arise even from pure rotational motion, introducing undesirable noise. To eliminate such disturbances, we initially fix a mounting point for each object and manually measure the directional offset $\vec{\bm{r}}$ from the center to that point using mesh processing software [9]. By leveraging recorded angular velocity $\vec{\bm{w}}$, the additional linear velocity stemming from rotation is $\vec{\bm{v}}=\vec{\bm{w}}\times\vec{\bm{r}}$, and we can calculate the excess linear acceleration: $\delta\bm{a}_{t}=\frac{\bm{v}_{t}-\bm{v}_{t-\Delta t}}{\Delta t}$. Finally, the normalized acceleration data can be attained by subtracting $\delta\bm{a}_{t}$ from the raw measurements.
@@
Furthermore, we simulate synthetic IMU data based on ground-truth object motion annotations of [4, 24, 102, 26]. In particular, to derive inertial acceleration data, we follow [93, 47, 67] to calculate the second-order difference of object translation:
@@
where $n=4$ is the smoothing factor to enhance the approximation to actual acceleration, and $\tau=\frac{1}{\text{fps}}$ represents the time interval between consecutive frames.